# GraphGSOcc: Semantic and Geometric Graph Transformer for 3D Gaussian Splating-based Occupancy Prediction

### v1

@@ 4 Experiments
In this section, we present the experimental setup and results obtained in this study, structured into three primary subsections for clarity. Section 4.1 details the dataset, evaluation metrics and training details employed in our experiments. In Section 4.2, we conduct a comparative analysis of our proposed DSOcc against several state-of-the-art 3D occupancy prediction approaches. Section 4.3 presents the ablation study, an essential part of our analysis that investigates the contribution of each component within our proposed framework.
@@ 4.1 Setting
Dataset. We performed occupancy analysis using the nuScenes dataset, which was captured by a data collection vehicle equipped with a multi-modal sensor array, including one LiDAR, five radars, and six cameras, ensuring full 360° environmental coverage. The dataset is partitioned into 700 training and 150 validation scenes, each lasting 20 seconds. High-precision ground truth annotations are provided at 0.5-second intervals, ensuring fine temporal resolution. Spatially, the dataset spans ±40 meters along the x- and y-axes and ranges from -1 to 5.4 meters in the z-axis. Occupancy labels are voxelized at a resolution of $0.4m\times 0.4m\times 0.4m$, covering 17 semantic categories as defined in SurroundOcc [15]. The perceptual data for each driving scene is annotated at 2 Hz, offering a comprehensive and detailed representation for occupancy analysis.
@@ 4.2.1 Quantitative analysis
When benchmarked against previous Gaussian-based techniques, GraphGSOcc demonstrates consistent superiority across multiple performance metrics. By integrating our proposed framework, the resulting model outperforms GaussianFormer by significant margins, achieving a 2.94% increase in IoU and a 1.83% improvement in mIoU. Beyond these accuracy gains, the integration also yields notable efficiency enhancements. Specifically, the inference latency is reduced by 132 ms, enabling faster processing times, while GPU memory usage is decreased by a substantial 2909 mb.
The performance improvements are not limited to GaussianFormer. When integrated with GaussianFormer-2, GraphGSOcc enables the model to achieve the lowest GPU memory consumption of just 2012 MB. As illustrated in Table 1, the best performance gains are observed when integrating GraphGSOcc into the GaussianWorld framework. These results collectively demonstrate the versatility and effectiveness of our approach across different Gaussian-based architectures, establishing GraphGSOcc as a powerful and practical solution for 3D occupancy prediction tasks.
@@ 4.2.2 Qualitative analysis
Our method is compared with high-performance benchmark, GaussianFormer, as illustrated in Figs. 6. In Fig. 6, we highlight the advantages of our proposed method compared to the other method with red circles. In the first row of Fig. 6, the visualization vividly demonstrates the efficacy of our proposed method in predicting the "driveable surface" regions that are otherwise obscured by much "cars". In the second and third rows of Figs. 6, GaussianFormer incorrectly predicted "car" and "bus" as "truck," respectively. These results indicate that the proposed method can effectively alleviate the semantic and geometric structure confusion caused by the overlap between different Gaussians. Ultimately, the method we proposed is capable of accurately predicting semantic categories and geometric structures.
Table 1: Performance comparison of 3D Occupancy Prediction methods on the nuScenes validation set. The latency and memory consumption for all methods are tested on one NVIDIA 4090 GPU with batch size one. The bold and underline indicate the best and second-best results, respectively. (TPVFormer* is supervised by dense occupancy annotations. GraphGSOcc** denotes ours method based on GaussianFormer. GraphGSOcc* denotes ours method based on GaussianFormer2. GraphGSOcc denotes our method based on GaussianWorld)
MonoScene CVPR’22 ResNet-101 1600 $\times$ 928 7.31 23.96 - -
BEVFormer ECCV’22 ResNet-101 1600 $\times$ 900 16.75 30.50 212 6651
TPVFormer CVPR’23 ResNet-101 1600 $\times$ 928 11.66 11.51 341 6926
TPVFormer* CVPR’23 ResNet-101 1600 $\times$ 928 17.10 30.86 341 6926
OccFormer ICCV’23 ResNet-101 1600 $\times$ 928 19.03 31.39 - -
SurroundOcc ICCV’23 ResNet-101 1600 $\times$ 900 20.30 31.49 312 5491
GraphGSOcc** - ResNet-101 1600 $\times$ 864 21.31 33.41 240 3320
GraphGSOcc* - ResNet-101 1600 $\times$ 864 22.65 34.68 232 2012
GraphGSOcc - ResNet-101 1600 $\times$ 864 24.10 35.21 305 6063
Table 2: Ablation on the Proposed DGGA Layer. Dyna. object denotes dynamic object. Stat. object denotes static object.
Methods mIoU IoU
All Dyna. object Stat. object All
baseline 18.41 16.73 20.09 29.23
w/ MLP 19.10 16.99 21.21 29.83
w/ GGA 19.83 17.21 22.45 31.98
w/ SGA 20.52 18.33 22.71 30.21
w/ DGGA 21.01 18.54 23.48 32.32
Table 3: Ablation on the MGA module.
Methods mIoU IoU Memory (MB)
Add 19.52 31.25 3080
Concat 19.48 31.10 3288
Conv Fusion 20.21 31.58 3250
Adaptive Fusion 21.31 33.41 3320
@@ 4.3.1 DGGA module
Table 4: Ablation on the parameters of Top-K and Top-M module.
Methods Top-K Top-M Memory (MB) mIoU IoU
A [100, 100, 100, 100] [100, 100, 100, 100] 3500 21. 22 33.22
B [200, 150, 100, 50] [200, 150, 100, 50] 3865 21.12 32.98
C [50, 25, 10, 5] [50, 25, 10, 5] 2010 19.51 31.21
D [100, 75, 50, 20] [100, 75, 50, 20] 3320 21.31 33.41
@@ 5 Conclusion
In conclusion, our proposed GraphGSOcc model marks a significant advancement in the realm of 3D semantic occupancy prediction within the context of vision-centric autonomous driving perception. By leveraging the potential of 3D Gaussian Splating (3DGS), GraphGSOcc effectively addresses the limitations of existing 3DGS-based methods, offering a more efficient and accurate solution for capturing the complex 3D structure of real-world scenes.
The Dual Gaussian Graph Attention (DGGA) mechanism and the Multi-scale Graph Attention (MGA) framework, as the core components of GraphGSOcc, play a crucial role in enhancing the model’s performance. The dynamic construction of geometric and semantic graphs through DGGA enables the model to adaptively optimize both boundary details and object-level topology, which effectively mitigates position drift and semantic ambiguities at object boundaries. Meanwhile, the MGA framework refines Gaussians hierarchically, optimizing both boundary details and small objects, which effectively mitigates position drift and small objects loss.
Our experimental results demonstrate the superiority of GraphGSOcc. Achieving a mIoU of 24.10% on Surroundocc and reducing GPU memory usage to 6.1 GB, our method not only improves the accuracy of 3D semantic occupancy prediction but also significantly enhances computational efficiency.

### v2

@@ 4 Experiments
In this section, we present the experimental setup and results obtained in this study, structured into three primary subsections for clarity. Section 4.1 details the dataset, evaluation metrics and training details employed in our experiments. In Section 4.2, we conduct a comparative analysis of our proposed GraphGSOcc against several state-of-the-art 3D occupancy prediction approaches. Section 4.3 presents the ablation study, an essential part of our analysis that investigates the contribution of each component within our proposed framework.
@@ 4.1 Setting
Dataset. We performed occupancy analysis using the SurroundOcc, Occ3D and OpenOcc based on nuScenes dataset, which was captured by a data collection vehicle equipped with a multi-modal sensor array, including one LiDAR, five radars, and six cameras, ensuring full $360^{\circ}$ environmental coverage. The dataset is partitioned into 700 training and 150 validation scenes, each lasting 20 seconds. High-precision ground truth annotations are provided at 0.5-second intervals, ensuring fine temporal resolution. Spatially, the dataset spans ±40 meters along the x- and y-axes and ranges from -1 to 5.4 meters in the z-axis. Occupancy labels are voxelized at a resolution of $0.4m\times 0.4m\times 0.4m$, covering 17, 17 and 16 semantic categories as defined in SurroundOcc [15], Occ3D [17] and OpenOcc [23], respectively. The perceptual data for each driving scene is annotated at 2 Hz, offering a comprehensive and detailed representation for occupancy analysis.
In addition to the nuScenes dataset, we also evaluate our approach on the KITTI-360 dataset [24] to validate the generality of our approach.
@@ 4.2.1 Quantitative analysis
When benchmarked against previous Gaussian-based techniques, GraphGSOcc demonstrates consistent superiority across multiple performance metrics. By integrating our proposed framework, the resulting model outperforms GaussianFormer by significant margins, achieving a 3.31% increase in IoU and a 4.38% improvement in mIoU. Beyond these accuracy gains, the integration also yields notable efficiency enhancements. Specifically, the inference latency is reduced by 92 ms, enabling faster processing times, while GPU memory usage is decreased by a substantial 2109 mb.
The performance improvements are not limited to GaussianFormer. When integrated with GaussianFormer-2, GraphGSOcc enables the model to achieve the lowest GPU memory consumption of just 2812 MB. As illustrated in Table 1, the best performance gains are observed when integrating GraphGSOcc into the GaussianWorld framework. These results collectively demonstrate the versatility and effectiveness of our approach across different Gaussian-based architectures, establishing GraphGSOcc as a powerful and practical solution for 3D occupancy prediction tasks.
Table 2 and Table 3 present the performance of our model on the Occ3D, OpenOcc, and KITTI-360 datasets, respectively. Our approach outperforms the previous state-of-the-art methods, demonstrating its strong generalization capability.
@@ 4.2.2 Qualitative analysis
Our method is compared with high-performance benchmark, GaussianFormer, as illustrated in Figs. 7 and Figs. 8. In Figs. 7 and Figs. 8, we highlight the advantages of our proposed method compared to the other method with red circles. In the first row of Fig. 7, the visualization vividly demonstrates the efficacy of our proposed method in predicting the "driveable surface" regions that are otherwise obscured by much "cars". In the second and third rows of Figs. 7, GaussianFormer incorrectly predicted "car" and "bus" as "truck," respectively. These results indicate that the proposed method can effectively alleviate the semantic and geometric structure confusion caused by the overlap between different Gaussians. Ultimately, the method we proposed is capable of accurately predicting semantic categories and geometric structures. Figs. 8 explicitly demonstrates that our model exhibits superior predictive performance and structural completeness for both dynamic objects and static scenes according to the red circles.
@@ 4.3.3 The parameters of Top-K and Top-M
4.3.4 DSDGA module
Table 7 demonstrates the effectiveness of the proposed DSDGA module. When this module is employed, the model exhibits better prediction performance for both moving objects and static scenes as a whole. The integration of the SCA leads to a significant improvement in the model’s predictive performance for static long-range contexts, while the adoption of the DCA results in a notable enhancement in the prediction of dynamic objects. This further validates that the decoupled interactive attention mechanism for dynamic and static objects can effectively facilitate the prediction of dynamic objects and static scenes.
Figs 12 presents the heatmap verifying the effectiveness of the DSDGA module. It can be observed from the heatmap that after incorporating DSDGA, the model demonstrates effective prediction for dynamic pedestrians, and successfully focuses features on moving pedestrians.
@@ 5 Conclusion
In conclusion, our proposed GraphGSOcc model marks a significant advancement in the realm of 3D semantic occupancy prediction within the context of vision-centric autonomous driving perception. By leveraging the potential of 3D Gaussian Splatting (3DGS), GraphGSOcc effectively addresses the limitations of existing 3DGS-based methods, offering a more efficient and accurate solution for capturing the complex 3D structure of real-world scenes.
The Dual Gaussian Graph Attention (DGGA) mechanism and the Multi-scale Graph Attention (MGA) framework, as the core components of GraphGSOcc, play a crucial role in enhancing the model’s performance. The dynamic construction of geometric and semantic graphs through DGGA enables the model to adaptively optimize both boundary details and object-level topology, which effectively mitigates position drift and semantic ambiguities at object boundaries. Meanwhile, the MGA framework refines Gaussians hierarchically, optimizing both boundary details and small objects, which effectively mitigates position drift and small objects loss. Additionally, the Dynamic-Static Decoupled Gaussian Attention (DSDGA) mechanism decouples dynamic and static objects, enabling the model to optimize the prediction for dynamic and static objects separately.
Our experimental results demonstrate the superiority of GraphGSOcc. Achieving a mIoU of 25.20% on Surroundocc and reducing GPU memory usage to 6.8 GB, our method not only improves the accuracy of 3D semantic occupancy prediction but also significantly enhances computational efficiency.